OpenStudy (anonymous):

The equation for the circle below is x2 + y2 = 64. What is the length of the circle's radius?

OpenStudy (anonymous):

\[x^2+y^2 = r^2\] is the form of the equation with r = radius

OpenStudy (anonymous):

is it 9?

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

No, \[r^2 = 64\] so what does r equal?

OpenStudy (whpalmer4):

no...what number * itself gives you 64?

OpenStudy (anonymous):

8

OpenStudy (anonymous):

Correct